OpenStudy (anonymous):

How do i solve √6(2+√6) Step-by-Step Plz.

OpenStudy (anonymous):

\[\sqrt{6}(2+\sqrt{6})=2\sqrt6+\sqrt6\cdot\sqrt6=2\sqrt6+\sqrt{6\cdot6}=2\sqrt6+\sqrt{36}=2\sqrt6+6=2(\sqrt6+3)\]

OpenStudy (anonymous):

last expression is not seen so I rewrite it \[2(\sqrt6+3)\] it's not necessary to do it but you can do it